OpenStudy (anonymous):

What is the maximum height, And when does it occur? h(t)=-16t^(2)+80t+3. Please help.

OpenStudy (anonymous):

the maximum height of this parabola is basically the vertex. so if you find the vertex, then there's your maximum height.

OpenStudy (anonymous):

Have you learned about derivatives yet?

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

Find the derivative of your function, and set it equal to zero, and solve for t. That will tell you when it occurred.

OpenStudy (anonymous):

So 0 = Derivative of function.

OpenStudy (anonymous):

okay thx guys :)

OpenStudy (anonymous):

If you stay on here, I can verify your work as you work through it.

OpenStudy (anonymous):

is it f(0)=83 or (0,80)

OpenStudy (anonymous):

Still solving that, but that doesn't seem correct.

OpenStudy (anonymous):

ill keep trying.

OpenStudy (anonymous):

Are you sure it's +80t and not -80t? From what I know, negative time doesn't exist.

OpenStudy (anonymous):

Sorry, the other way around.

OpenStudy (anonymous):

the problem says h(t)=-16t^(2)+80t+3

OpenStudy (anonymous):

Aaaah! I see. When you posted your problem, the negative sign was cut off so all I saw was 16t^2

OpenStudy (anonymous):

Alright, I solved it.

OpenStudy (anonymous):

Where do you need help?

OpenStudy (anonymous):

If you're just completely lost I can walk you through it.

OpenStudy (anonymous):

that would help ALOT.

OpenStudy (anonymous):

Sounds good! Okay, so your function is f(t) = -16t^2 + 80t + 3 The derivative of that function is (df/dt) = -32t + 80 Does that make sense?

OpenStudy (anonymous):

yess.

OpenStudy (anonymous):

The function reaches its maximum height when the derivative, or the slope, is equal to zero. It's kind of like when you throw a ball in the air. The moment the ball reaches the top of its arc, that's when the slope is zero. So what you do is set the derivative to zero ... -32t + 80 = (df/dt) -32t + 80 = 0 Given that easy formula, solve for t and let me know what you get.

OpenStudy (anonymous):

It's just algebra at this point.

OpenStudy (anonymous):

does t= 5/2?

OpenStudy (anonymous):

Yeah! Awesome! Now, you plug that into h(t) and the number you get is your height.

OpenStudy (anonymous):

And then you're done! :D

OpenStudy (anonymous):

THANKS SO MUCH!

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

I HAve Another One or You.

OpenStudy (anonymous):

*for